OpenStudy (anonymous):

If the formula for a compound is represented by X3Y2 and the charge on the Y ion is -3, what is the charge in the X ion?

OpenStudy (rushwr):

+2

OpenStudy (anonymous):

IT would be +2 correct?

OpenStudy (rushwr):

yp !

OpenStudy (rushwr):

X3Y2 Charge from Y = -3*2 = -6 The charge from 3 molecules of X = +6 The charge of one X molecule = +6/3 = +2